40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Nov 7 10:52:57 2008, steve, Update, PEM, AC is back on acf.jpgacc_mc1.jpgtest.jpg
    Reply  Fri Nov 7 15:58:10 2008, rana, Update, PEM, AC is back on mc-acc.pdfmc-acc-quad.pdf
       Reply  Mon Nov 10 19:39:40 2008, steve, Update, PEM, thump noise from particle counter partcfm.jpg
          Reply  Tue Nov 12 23:05:06 2013, rana, Update, PEM, thump noise from particle counter 
             Reply  Wed Nov 13 11:30:07 2013, Steve, Update, PEM, particle counter MOVED Met1Moved.JPG
Message ID: 9369     Entry time: Tue Nov 12 23:05:06 2013     In reply to: 1129     Reply to this: 9374
Author: rana 
Type: Update 
Category: PEM 
Subject: thump noise from particle counter 

Quote:
The particle counter is back on the IOOC location on a piece of FOAM
It needs this isolation, so when the pump is running, it's not shaking things.
The counter was counting for 6 sec and it was on holding for 20 mins.
Now I set the counter for 20 sec so it is easy to recognise it's signal and it holds for 2min only.
This will set the alarm handler in action.

Atm1: 40 mins plot
PEM-ACC_MC2_x,y,z up to 13 mins: pcounter at MC2 table, clamped, counting for 20s and holds for 2 mins
PEM-ACC_MC2_x,y,z from 13 to 26 mins: pcounter at MC2 table, not clamped, seated on 2" foam, counting 20s and holds for 2 mins
PEM-ACC_MC1_x,y,z from 26 to 40 mins: pcounter at MC1_IOOC location, not clamped, seated on 2" foam, counting 20s and holds for 2 mins

Rana won the bet

This noise source from 2008 is back. Somehow the particle counter snuck back up to the top of the BS chamber. Bad, noisy particle counter. Let's move it away from out sensitive optics and put it on the wall next to the router or maybe on some toolbox. Not on optics tables, chambers, or near them.
ELOG V3.1.3-